OpenStudy (qaisersatti):

thanks god finally its done 100 percent with the help of all of u friends..if someone needs help message me i will be replying tomorrow morning ..

OpenStudy (anonymous):

the formula of RTH plz

OpenStudy (anonymous):

what abut q6 and q4

OpenStudy (anonymous):

q4 1-(R*b*vIN)/(1+R*b) 2-(VS*(1+R*b)-vIN)/(1+R*b)

OpenStudy (anonymous):

Formulas for Q3? "For the last two parts assume I = 2A" Also, formulas for Q6 please :)

OpenStudy (anonymous):

plzz plzz Q3 and Q6

OpenStudy (anonymous):

q6 Vin=45V, R1=1.5Ω, R2=2Ω, R3=6Ω, A=2 and B=0.5 1-i1=5 2-vo=7.5 3-vth=7.5

OpenStudy (anonymous):

formulae

OpenStudy (anonymous):

plzz post formulae

OpenStudy (anonymous):

Thx EngSara

OpenStudy (anonymous):

q6 the formula 1-Vo* (R3-(B*R1*R3)+R1+R2+A*R2)/(R3-B*R1*R3)=VIN 2-i1=V0/(R3-B*R1*R3) 3- vo = vth

OpenStudy (anonymous):

Someone with Rth formula , would be appreciated !

OpenStudy (anonymous):

RTH???

OpenStudy (anonymous):

@ali_amjad , Brother , it is the last part of question 6

OpenStudy (anonymous):

Anyone has the formulas of the Q3?

OpenStudy (anonymous):

i don't know the Rth

OpenStudy (anonymous):

@EngSara , same is the case here :-/

OpenStudy (anonymous):

what about Q3

OpenStudy (anonymous):

q3 if I=10 vo=2 if I=-3 vo= -1 if I =1 vo=0.333 v1=0.665

OpenStudy (anonymous):

plzz the formula aor hint for Q3

OpenStudy (anonymous):

the resistance of zener =1

OpenStudy (anonymous):

Thx EngSara Again

OpenStudy (anonymous):

you are welcome :)

OpenStudy (anonymous):

i am trying Q3

OpenStudy (anonymous):

@EngSara , how did you find out that if I = 10 , then vo = 2 ????

OpenStudy (anonymous):

@EngSara not getting it in Q3

OpenStudy (anonymous):

@chaand.i i dont thimk we sholud think zener as R=1 because if so then v0=7.5 . .??

OpenStudy (anonymous):

@EngSara ????

OpenStudy (benzo):

engsara please help me with q3 what the formula or answer I 9a then vo =...I is -3 vo=..Iis 4 then v1= and vo=

OpenStudy (anonymous):

chaand.i its to 3,3V but the zener cuts it down to 2V

OpenStudy (anonymous):

Q3 for I=2A? What's the formula? :3

OpenStudy (anonymous):

@Zozi what the formula for Q3

OpenStudy (anonymous):

\[(2ohm/3ohm)*I\] for v1 and for v2 its half, but over two volts the zener cuts dovn the voltage to 2V

OpenStudy (anonymous):

RTH pleasee

OpenStudy (benzo):

engsara please help......q6 I1, Vo, Vth, Rth formula pleaseeee?

OpenStudy (anonymous):

Q6 Rth please, Vin=18V, R1=1.5Ω, R2=3Ω, R3=4Ω, A=1.5 and B=0.5.

OpenStudy (anonymous):

i got it My Rth= 3.6

OpenStudy (anonymous):

Q3: p1 : zener voltage sholud be less than the given current so vo=2V which is zener voltage Q3: P2 : I1=( 0.66 * I ) . . . also find I2 = (0.33 * I) .. . Vo= I2 Q3: P3 : (a) I1=( 0.66 * I ) . . . also find I2 = (0.33 * I) .. . Vo= I2 . . (b) v1 = I1

OpenStudy (anonymous):

@JoãoVitorMC how you got it?? Rth

OpenStudy (anonymous):

I got my Rth here http://openstudy.com/study#/updates/5089b78de4b0d596c4603b9e

OpenStudy (anonymous):

@chaand.i the volt must be 3.33 but when see the fig of zener , the max volt is 2

OpenStudy (anonymous):

@JoãoVitorMC : if you consider zener as R=1phm resistor then this question would be much much simpler

OpenStudy (anonymous):

R=1ohm

OpenStudy (anonymous):

yes that's what i do, but this happens only then the voltage saw by the element is in the range -+2

OpenStudy (anonymous):

yup because in the first part the current source is too big and then zener limits the voltage

OpenStudy (anonymous):

Rth ???/

OpenStudy (anonymous):

thank you @EngSara and @Zozi for help of Q3

OpenStudy (anonymous):

RTH= Vth / ISC Isc= Current on short circuit How can I find the ISC?